# Fair and Truthful Allocations Under Leveled Valuations

### v1

@@ 1 Introduction
The maximin share guarantee (MMS) is a compelling and well-studied notion proposed by Budish (2010) which can be seen as a natural extension of the well-known cut-and-choose protocol. The goal here is to allocate a bundle to each agent, ensuring that its value is at least as high as her maximin share (or a large fraction thereof). Intuitively, the maximin share corresponds to the maximum value an agent can get after proposing an initial allocation and keeping the least desirable bundle for herself. The study of MMS allocations and approximations thereof (mostly under additive utilities) has received significant attention over the last few years Kurokawa et al. (2018), Amanatidis et al. (2017b), Barman and Krishnamurthy (2020), Garg and Taki (2021).
In sharp contrast to weaker relaxations, the existence of envy-freeness up to any good (EFX) allocations Caragiannis et al. (2019) is yet to be settled beyond a handful of simple settings Plaut and Roughgarden (2020), Chaudhury et al. (2020). A reasonable approach is to focus on restricted valuation spaces and approximations.
In this work, we focus on a different class of valuations, namely leveled valuations. They capture a simple economic behavior in which any set of items is strictly preferred to any other set of smaller cardinality, regardless of the specific items included; among bundles of equal size any preference ordering is feasible. [^1] Leveled preferences can be perceived as a way of prioritizing quantity over quality, thus giving rise to the definition of quantity-monotonic preferences used by Pápai (2000). Babaioff et al. (2021b) study leveled preferences in the context of competitive equilibrium in markets, while Gafni et al. (2023) study EFX allocations under leveled preferences for chores. We also note that leveled preferences might arise in the context of web advertising where some competitors aim to increase their total coverage rather than focusing on specific advertising spots in the ad space allocation. Lastly, leveled preferences might appear under budget constraints: when resources are scarce, agents facing budget constraints might prioritize quantity to ensure that they have enough items to meet their needs.
@@ 1.1 Our Contributions
In light of this negative result, and in order to provide a more fine-grained analysis of approximate MMS under leveled valuations, we focus on important valuation subclasses such as submodular leveled (Section 3.1), fractionally subadditive (Section 3.2) and subadditive leveled functions (Section 3.3). Submodular, XOS, and subadditive valuations capture complex combinatorial preferences and have been studied extensively in microeconomics and algorithmic game theory Lehmann et al. (2006), Feige (2009), Christodoulou et al. (2016), Feldman et al. (2020).
@@
In Section 4 we shift our attention to EFX allocations. We prove the existence of exact EFX allocations under (ordinal or cardinal) leveled preferences via a simple protocol. In particular, under strict leveled preferences our mechanism achieves a set of natural properties in addition to EFX, overcoming the barriers of existing impossibility results on the interplay between fairness, efficiency, and incentive compatibility. Our results complement the previous work of Gafni et al. which showed the existence of EFX allocations under the special case of additive leveled valuations. [^2] We note that the model studied in Gafni et al. (2023) is different than ours. They consider fair allocations for goods with copies and develop a duality framework between chores and goods with copies. Without copies, their proposed fairness concept coincides with the standard notion of EFX.
@@
Finally, we explore the interplay of fairness with truthfulness for the class of leveled valuations. We are able to show that constant factor MMS approximations can be achieved via truthful mechanisms under leveled valuations. This is in contrast to (non-leveled) valuations where non-constant upper bounds persist even for the case of additive valuations Amanatidis et al. (2017a). Additionally, we show an algorithmic characterization of mechanisms satisfying truthfulness, Pareto optimality, non-bossiness, neutrality, and EFX.
@@ 1.2 Related Work
In this section we discuss prior works closely related to EFX, MMS, and leveled valuations. The growing literature on fair allocation of indivisible items is too extensive to cover on this paper and thus, we refer the interested reader to the recent survey of Amanatidis et al. for a detailed overview.
@@
EFX. The concept of EFX poses several challenges leading to important open problems in the area. Plaut and Roughgarden (2020) showed that when agents exhibit identical valuations or share the same ordinal ranking of goods, EFX is guaranteed to exist. In a breakthrough result, Chaudhury et al. (2020) showed that complete, exact EFX allocations always exist for three agents under additive utilities, later improved to more general valuations Berger et al. (2022), Akrami et al. (2023a). Other works considered a limited number of items or agents sharing one of two possible valuation types Amanatidis et al. (2020), Mahara (2023, 2024), while a major line of work focused on restrictions on the valuation space, prompting a series of works on dichotomous valuations Amanatidis et al. (2021), Babaioff et al. (2021a), Benabbou et al. (2021), Halpern et al. (2020), lexicographic preferences Hosseini et al. (2021), graph instances Christodoulou et al. (2023), and allocations of multisets Gorantla et al. (2023). Recent work has shown that EFX allocations might not exist under general monotone valuations when the items to be allocated are chores Christoforidis and Santorinaios (2024).
@@
MMS. The existence of maximin fair allocations is not guaranteed for additive utilities Kurokawa et al. (2018). A series of works established strong approximation guarantees reaching a factor of $3/4+\mathcal{O}(1/n)$ Kurokawa et al. (2018), Amanatidis et al. (2017b), Barman and Krishnamurthy (2020), Ghodsi et al. (2021), Garg and Taki (2021), Akrami et al. (2023b). Very recently, Akrami and Garg (2024) broke the barrier of $3/4$. Considerable effort has been shown to the case of four or fewer agents Amanatidis et al. (2017b), Gourvès and Monnot (2019), Ghodsi et al. (2021), Feige and Norkin (2022) or settings with few items Feige et al. (2021), Hummel (2023). The current impossibility result for additive valuations stands at $1-1/n^{4}$ due to Feige et al. (2021). Moreover, the study of maximin fair allocations beyond the additive domain has been of growing interest. Barman and Krishnamurthy (2020), Uziahu and Feige (2023) and Ghodsi et al. (2022), Seddighin and Seddighin (2024) provided a series of results for complement-free valuations. We briefly summarize the known results in Table 1.
@@
Leveled Preferences. Typical examples of such preferences mentioned in the literature include the distribution of cabinet seats among parties forming a coalition government or the allocation of offices among different departments at universities. Pápai (2000) characterized the set of strategyproof mechanisms under standard economic assumptions in assignment problems. Gafni et al. (2023) prove the existence of EFX allocations under leveled additive valuations, while Babaioff et al. (2021a) consider $\epsilon$-leveled valuations (additive valuation functions that are $\epsilon$-dichotomous). Oh et al. (2021) show that there exists a deterministic algorithm that computes an EF1 allocation using $\mathcal{O}(n^{2})$ queries under leveled valuations. Despite their simplicity, leveled valuations have played an important role in constructing counterexamples; we note that the first negative results for both notions (MMS non-existence and exponential query complexity for EFX) actually employ leveled valuations Kurokawa et al. (2018), Plaut and Roughgarden (2020). The best currently known MMS bounds for more general classes are also (essentially) leveled.
@@ 2 Model and Preliminaries
Efficiency. A common notion of economic efficiency is Pareto optimality. Roughly speaking, under a Pareto efficient allocation, no individual can become strictly better off without hurting someone else’s value. There exist instances where EFX cannot be satisfied in conjunction with Pareto optimality, even for 2 players Plaut and Roughgarden (2020).
@@ 3 Maximin share (MMS)
In this section, we focus on approximate MMS and show general positive and negative results under leveled valuations. First, we focus on instances with few items; we show that every instance with $m<2n$ admits an exact MMS allocation under general leveled valuations. [^3] Similar settings with a small number of items have been explored in a surge of works for additive utilities Bouveret and Lemaître (2016), Kurokawa et al. (2018), Feige et al. (2021), Hummel (2023). Exceeding this threshold may lead to instances where the approximation ratio becomes unbounded. Subsequently, we proceed to a more refined analysis of the concept for valuations that are both leveled and submodular. We show a general positive result, establishing the existence of $2/3$-MMS allocations under submodular leveled valuations (Theorem 1). We then switch to subadditive and fractionally subadditive agents and show tight lower and upper bounds on the approximation of the maximin share (Section 3.2 and Section 3.3). We now state our proposition concerning scenarios where the number of items is limited in relation to the number of agents, specifically when $m<2n$.
@@
For $n\geq m$ verifying the existence of MMS allocations is straightforward. In the scenario where $n<m<2n$, it is clear that the maximin value for each agent is achieved via singletons. Suppose that we have $m=n+r$ with $r$ a positive integer smaller than $n$. Select an arbitrary picking order $\sigma$, a permutation of $[n]$. We let the first $n-r$ agents choose their preferred item according to $\sigma$. Subsequently, we allocate the rest of the items in pairs to the remaining agents in an arbitrary manner. Clearly, agents receiving two items guarantee their full MMS due to leveled valuations. In the case of an agent $i$ receiving a single item, it holds that all the agents that came beforehand in the picking order have received a single item too. Removing an arbitrary item $g$ and an arbitrary agent $j$ does not decrease the MMS value for the remaining agents. Such reducibility lemmas have been used several times in the literature Bouveret and Lemaître (2016), Amanatidis et al. (2023). ∎
@@
Consider an instance with $n$ agents and $m=2n$ goods $\{g_{1},\dots,g_{m}\}$. For all agents $i\in[n-1]$ we have that $v_{i}(S)=\epsilon$ for all $S$, such that $\lvert S\rvert=1$. Furthermore, $v_{i}(\{g_{2j-1},g_{2j}\})=1,\forall j\in[n],\forall i\in[n]$, while any other pair of items is valued at $\epsilon+\delta$, with $\epsilon,\delta$ small positive constants. For agent $n$ we have $v_{n}(S)=\epsilon$ for all $S$, $\lvert S\rvert=1$, and $v_{n}(\{S^{\prime}\})=1$ for all bundles $S^{\prime}\in\mathcal{S}=(\{g_{1},g_{2n}\},\{g_{2},g_{3}\},\dots,\{g_{2n-2},g_{2n-1}\})$; the rest of the pairs are also valued at $\epsilon+\delta$. Lastly, $v_{i}(S)=\lvert S\rvert+\epsilon$ for all $\lvert S\rvert>2$. One can easily check that the valuations described above are leveled. The proof follows the structure of the proof of Corollary 2 due to Ghodsi et al. (2022), where all but one agent share the same valuation function. Agent $1$’s maximin share is equal to $1$; same holds true for agent $2$. We observe that one must receive a low-valued bundle. It is easy to see that we can make $\epsilon$ infinitesimally small while keeping the valuations leveled, thus making any approximation guarantee arbitrarily bad. Hence, we shift our attention to subclasses of leveled valuation functions, such as submodular, fractionally subadditive, and subadditive leveled valuations. Our results are summarized in Table 1.
@@
Table 1: Best known approximations of MMS with submodular agents are due to Ghodsi et al. (2022), Uziahu and Feige (2023), Kulkarni et al. (2023), Seddighin and Seddighin (2024). The results on XOS valuations are due to Ghodsi et al. (2022) and Akrami et al. (2023c), while bounds on subadditive valuations come from Ghodsi et al. (2022) and Seddighin and Seddighin (2024). Our results appear in bold. The upper bounds noted with $(^{*})$ are not explicitly stated in previous works, but one can tweak known counterexamples to also work for leveled valuations.
@@ 3.1.1 Lower Bound for Two Submodular Agents
3.2 Fractionally Subadditive Leveled Valuations
We exhibit tight bounds for the case of fractionally subadditive (XOS) leveled valuations. The existence result follows from Section 3.3. The non-existence (impossibility) construction relies on a family of additive functions that replicate the $1/2$ bound given in Section 3.3.
@@ 3.3 Subadditive Leveled Valuations
Fix an agent $i$ and two disjoint subsets $S,T$ of $M$ with $|S|=|T|=\lfloor m/n\rfloor$. If $v_{i}(S)<1/2$, then $v_{i}(T)>1/2$.
@@ 3.4 Upper Bounds
For subadditive valuations, one can adapt the construction given in Ghodsi et al. (2022).
@@
There exists an instance of the fair allocation problem with two submodular leveled agents in which no allocation is strictly better than 2/3-MMS Kulkarni et al. (2023).
@@
For any $n\geq 2$ there exists an instance of the fair allocation problem with n subadditive leveled agents where no allocation yields more than $1/2$-MMS to the agents Ghodsi et al. (2022).
@@ 5 Truthfulness
In this section we study the mechanism design aspect of fairly allocating goods among agents. In Section 5.1 we show that constant MMS approximations are possible via a rich class of truthful mechanisms. In Section 5.2 we show a characterization of truthful, EFX, PO, non-bossy, and neutral mechanisms. Under leveled valuations, we observe that strategyproofness can be achieved in conjunction with strong fairness guarantees. Amanatidis et al. introduce the notion of controlling items; we say that a player controls an item if reporting certain values guarantees her this item. Control is the major hurdle that makes fairness and truthfulness incompatible under additive valuations. We note that this is not applicable to our case, since control of singletons [^6] Control of singletons implies that whenever an agent strongly desires an item, i.e. prefers it more than all the other items combined, she gets it. (or pairs) is not possible under leveled valuations.
The mechanism described in Section 4 falls under the broad class of serial dictatorships with quotas (see Definition 5).

### v2

@@ 1 Introduction
The maximin share guarantee (MMS) is a compelling and well-studied notion proposed by Budish (2011) which can be seen as a natural extension of the well-known cut-and-choose protocol. The goal here is to allocate a bundle to each agent, ensuring that its value is at least as high as her maximin share (or a large fraction thereof). Intuitively, the maximin share corresponds to the maximum value an agent can get after proposing an initial allocation and keeping the least desirable bundle for herself. Such allocations might not exist for additive valuations (Kurokawa et al., 2018). The study of MMS allocations and approximations thereof (mostly under additive utilities) has received significant attention over the last few years (Kurokawa et al., 2018, Amanatidis et al., 2017b, Barman and Krishnamurthy, 2020, Garg and Taki, 2021).
In sharp contrast to weaker relaxations, the existence of envy-freeness up to any good (EFX) allocations (Caragiannis et al., 2019) is yet to be settled beyond a handful of simple settings (Plaut and Roughgarden, 2020, Chaudhury et al., 2024). A reasonable approach is to focus on restricted valuation spaces and approximations.
In this work, we focus on a specific class of valuations, namely leveled valuations. They capture a simple economic behavior in which any set of items is strictly preferred to any other set of smaller cardinality, regardless of the specific items included; among bundles of equal size any preference ordering is feasible. [^1] Leveled preferences can be perceived as a way of prioritizing quantity over quality, thus giving rise to the definition of quantity-monotonic preferences used by Pápai (2000). Babaioff et al. (2021b) study leveled preferences in the context of competitive equilibrium in markets, while Gafni et al. (2023) study EFX allocations under leveled preferences for chores. We also note that leveled preferences might arise in the context of web advertising where some competitors aim to increase their total coverage rather than focusing on specific advertising spots in the ad space allocation. Lastly, leveled preferences might appear under budget constraints: when resources are scarce, agents facing budget constraints might prioritize quantity to ensure that they have enough items to meet their needs.
@@ 1.1 Our Contributions
In light of this negative result, and in order to provide a more fine-grained analysis of approximate MMS under leveled valuations, we focus on important valuation subclasses such as submodular leveled (Section 3.1), fractionally subadditive (XOS) and subadditive leveled functions (Section 3.2). Submodular, XOS, and subadditive valuations capture complex combinatorial preferences and have been studied extensively in microeconomics and algorithmic game theory (Lehmann et al., 2006, Feige, 2009, Christodoulou et al., 2016, Feldman et al., 2020).
@@
In Section 4 we shift our attention to EFX allocations. We prove the existence of exact EFX allocations under leveled valuations via a simple protocol. Moreover, our algorithm works even when only ordinal information is given as input. In particular, under strict leveled preferences our mechanism achieves a set of natural properties in addition to EFX, overcoming the barriers of existing impossibility results on the interplay between fairness, efficiency, and incentive compatibility. Our results complement the previous work of Gafni et al. (2023) which showed the existence of EFX allocations under the special case of additive leveled valuations. [^2] We note that the model studied in (Gafni et al., 2023) is different than ours. They consider fair allocations for goods with copies and develop a duality framework between chores and goods with copies. Without copies, their proposed fairness concept coincides with the standard notion of EFX.
@@
Finally, we explore the interplay of fairness with truthfulness for the class of leveled valuations. We are able to show that constant factor MMS approximations can be achieved via truthful mechanisms under leveled valuations. This is in contrast to (non-leveled) valuations where non-constant upper bounds persist even for the case of additive valuations (Amanatidis et al., 2017a). Additionally, we show an algorithmic characterization of mechanisms satisfying truthfulness, Pareto optimality, non-bossiness, neutrality, and EFX.
We note that our focus is explicitly on the existence of such notions, rather than on their computational aspects.
@@ 1.2 Related Work
In this section we discuss prior works closely related to EFX, MMS, and leveled valuations. The growing literature on fair allocation of indivisible items is too extensive to cover on this paper and thus, we refer the interested reader to the recent survey of Amanatidis et al. (2023) for a detailed overview.
@@
EFX. The concept of EFX poses several challenges leading to important open problems in the area. Plaut and Roughgarden (2020) showed that when agents exhibit identical valuations or share the same ordinal ranking of goods, EFX is guaranteed to exist. In a breakthrough result, Chaudhury et al. (2024) showed that complete, exact EFX allocations always exist for three agents under additive utilities, later improved to more general valuations (Berger et al., 2022, Akrami et al., 2023a). Other works considered a limited number of items or agents sharing one of two possible valuation types (Amanatidis et al., 2020, Mahara, 2023, 2024), while a major line of work focused on restrictions on the valuation space, prompting a series of works on dichotomous valuations (Amanatidis et al., 2021, Babaioff et al., 2021a, Benabbou et al., 2021, Halpern et al., 2020), lexicographic preferences (Hosseini et al., 2021), graph instances (Christodoulou et al., 2023), and allocations of multisets (Gorantla et al., 2023). Recent work has shown that EFX allocations might not exist under general monotone valuations when the items to be allocated are chores (Christoforidis and Santorinaios, 2024).
@@
MMS. The existence of maximin fair allocations is not guaranteed for additive utilities (Kurokawa et al., 2018). A series of works established strong approximation guarantees reaching a factor of $3/4+\mathcal{O}(1/n)$ (Kurokawa et al., 2018, Amanatidis et al., 2017b, Barman and Krishnamurthy, 2020, Ghodsi et al., 2021, Garg and Taki, 2021, Akrami et al., 2023b). Very recently, Akrami and Garg (2024) broke the barrier of $3/4$. Considerable effort has been shown to the case of four or fewer agents (Amanatidis et al., 2017b, Gourvès and Monnot, 2019, Ghodsi et al., 2021, Feige and Norkin, 2022) or settings with few items (Feige et al., 2021, Hummel, 2023). The current impossibility result for additive valuations stands at $1-1/n^{4}$ due to Feige et al. (2021). Moreover, the study of maximin fair allocations beyond the additive domain has been of growing interest. Barman and Krishnamurthy (2020), Uziahu and Feige (2023) and Ghodsi et al. (2022), Seddighin and Seddighin (2024) provided a series of results for complement-free valuations. We briefly summarize the known results in Table 1.
@@
Leveled Preferences. Typical examples of such preferences mentioned in the literature include the distribution of cabinet seats among parties forming a coalition government or the allocation of offices among different departments at universities. Pápai (2000) characterized the set of strategyproof mechanisms under standard economic assumptions in assignment problems. Gafni et al. (2023) prove the existence of EFX allocations under leveled additive valuations, while Babaioff et al. (2021a) consider $\epsilon$-leveled valuations (additive valuation functions that are $\epsilon$-dichotomous). Oh et al. (2021) show that there exists a deterministic algorithm that computes an EF1 allocation using $\mathcal{O}(n^{2})$ queries under leveled valuations. Despite their simplicity, leveled valuations have played an important role in constructing counterexamples; we note that the first negative results for both notions (non-existence of MMS allocations and exponential query complexity for EFX) actually employ leveled valuations (Kurokawa et al., 2018, Plaut and Roughgarden, 2020). The best currently known MMS bounds for more general classes are also (essentially) leveled.
@@ 3 Maximin share (MMS)
In this section, we focus on approximate MMS and show general positive and negative results under leveled valuations. First, we focus on instances with few items; we show that every instance with $m<2n$ admits an exact MMS allocation under general leveled valuations. [^3] Similar settings with a small number of items have been explored in a surge of works for additive utilities (Bouveret and Lemaître, 2016, Kurokawa et al., 2018, Feige et al., 2021, Hummel, 2023). Exceeding this threshold may lead to instances where the approximation ratio becomes unbounded. Subsequently, we proceed to a more refined analysis of the concept for valuations that are both leveled and submodular. We show a general positive result, establishing the existence of $2/3$-MMS allocations under submodular leveled valuations (Theorem 1). We then switch to subadditive and fractionally subadditive agents and show tight lower and upper bounds on the approximation of the maximin share (Section 3.2 and Section 3.3).
We now state our proposition concerning scenarios where the number of items is limited in relation to the number of agents, specifically when $m<2n$.
@@
For $n\geq m$ verifying the existence of MMS allocations is straightforward. Let’s consider the case of $m=n+r$ for some positive $r<n$. Note that in that case, due to leveled valuations, the MMS partition of each agent consists of $n-r$ singletons and $r$ pairs, and the MMS value corresponds to the value of the $n-r$-th best singleton. Let $\sigma$ be some arbitrary permutation of $[n]$. We let the first $n-r$ agents w.r.t $\sigma$ to choose their preferred item. Subsequently, we allocate the rest of the items in pairs to the remaining agents in an arbitrary manner. Clearly, agents receiving two items guarantee their full MMS due to leveled valuations. In the case of an agent $i$ receiving a single item, it holds that there are are at most $n-r-1$ agents that precede her in $\sigma$ each getting at most one item. Therefore $i$ can still guarantee her MMS value. ∎
@@
Consider an instance with $n$ agents and $m=2n$ goods $\{g_{1},\dots,g_{m}\}$. For all agents $i\in[n-1]$ we have that $v_{i}(S)=\epsilon$ for all $S$, such that $\lvert S\rvert=1$. Furthermore, $v_{i}(\{g_{2j-1},g_{2j}\})=1,\forall j\in[n],\forall i\in[n-1]$, while any other pair of items is valued at $\epsilon+\delta$, with $\epsilon,\delta$ small positive constants. For agent $n$ we have $v_{n}(S)=\epsilon$ for all $S$, $\lvert S\rvert=1$, and $v_{n}(S^{\prime})=1$ for all bundles $S^{\prime}\in\mathcal{S}=(\{g_{1},g_{2n}\},\{g_{2},g_{3}\},\dots,\{g_{2n-2},g_{2n-1}\})$; the rest of the pairs are also valued at $\epsilon+\delta$. Lastly, $v_{i}(S)=\lvert S\rvert+\epsilon$ for all $\lvert S\rvert>2$. One can easily check that the valuations described above are leveled. The proof follows the structure of the proof of Corollary 3 due to Ghodsi et al. (2022), where all but one agent share the same valuation function. The maximin share of each agent is equal to $1$. We observe that one must receive a low-valued bundle. It is easy to see that we can make $\epsilon$ infinitesimally small while keeping the valuations leveled, thus making any approximation guarantee arbitrarily bad.
Hence, we shift our attention to subclasses of leveled valuation functions, such as submodular, fractionally subadditive, and subadditive leveled valuations. Our results are summarized in Table 1.
@@
Table 1: Best known approximations of MMS with submodular agents are due to Ghodsi et al. (2022), Uziahu and Feige (2023), Kulkarni et al. (2023), Seddighin and Seddighin (2024). The results on XOS valuations are due to Ghodsi et al. (2022) and Akrami et al. (2023c), while bounds on subadditive valuations come from (Ghodsi et al., 2022) and (Seddighin and Seddighin, 2024). Our results appear in bold. The upper bounds noted with $(^{*})$ are not explicitly stated in previous works, but one can tweak known counterexamples to also work for leveled valuations.
@@ 3.2 Subadditive Leveled Valuations
Fix an agent $i$ and two disjoint subsets $S,T$ of $M$ with $|S|=|T|=\lfloor m/n\rfloor$ and let $m\geq n$. If $v_{i}(S)<1/2$, then $v_{i}(T)>1/2$.
@@ 3.3 Upper Bounds
In this section, we provide impossibility results for approximating the MMS value for submodular (Corollary 1), XOS (Corollary 2) and subadditive (Corollary 3) leveled valuations which are simple adaptations of known impossibility constructions. In fact, the only necessary modification in their constructions, is to adjust the valuations to be strictly increasing in the number of items in order to make them leveled valuations.
Kulkarni et al. (2023) showcased a construction that gives an impossibility result on the existence of better than $2/3$-MMS allocations for two submodular agents. In fact, they show that the result holds even for a smaller subclass of submodular valuations, namely OXS functions. One can easily adapt their construction to make it work for leveled valuations (Corollary 1).
@@
For subadditive valuations, one can adapt the construction given in (Ghodsi et al., 2022).
@@
There exists an instance of the fair allocation problem with two submodular leveled agents in which no allocation is strictly better than 2/3-MMS (Kulkarni et al., 2023).
@@
For any $n\geq 2$ there exists an instance of the fair allocation problem with $n$ XOS leveled agents where no allocation yields more than $1/2$-MMS to the agents (Ghodsi et al., 2022).
Corollary 3.
For any $n\geq 2$ there exists an instance of the fair allocation problem with $n$ subadditive leveled agents where no allocation yields more than $1/2$-MMS to the agents (Ghodsi et al., 2022).
@@ 5 Truthfulness
In this section we study the mechanism design aspect of fairly allocating goods among agents. In Section 5.1 we show that constant MMS approximations are possible via truthful mechanisms, in contrast to non-leveled valuations. Specifically, we show that any balanced allocation, which allocates bundles of size $\lfloor m/n\rfloor$, and $\lceil m/n\rceil$, achieves a $\frac{k-1}{k}$-MMS approximation under additive leveled valuations for $k\geq 2$, where $k=\lfloor m/n\rfloor$. Furthermore, we show that this bound is tight for balanced allocations. The algorithm described in Section 4 satisfies truthfulness and outputs a balanced allocation, thus, simultaneously guarantees EFX, $\frac{k-1}{k}$-MMS, and truthfulness for additive leveled valuations. [^6] An earlier version of our work included the claim that the allocation obtained by Algorithm 1 is EFX and $\max\{2/3,(k-1)/k\}$-MMS, albeit with an incorrect proof for the case of $2/3$. We thank Mahyar Afshinmehr, Mehrafarin Kazemi, and Kurt Mehlhorn for pointing out a bug in our proof. In fact, they provide an improved bound, showing that the allocation satisfies EFX and $\frac{k}{k+1}$-MMS (Afshinmehr et al., 2024). Then, we show that an $\frac{1}{2}$-MMS approximation can be achieved for subadditive leveled valuations via a truthful mechanism.
In Section 5.2 we show a characterization of truthful, EFX, PO, non-bossy, and neutral mechanisms under strict ordinal leveled preferences.
5.1 Maximin Share Guarantee
In the case of additive agents, it is known that an EFX allocation is also a $2/3$-MMS allocation for $n\in\{2,3\}$, due to Amanatidis et al. (2018). The approximation degrades to $4/7=0.5714$ as the number of agents grows. We proceed to show that balanced allocations achieve a good MMS approximation in general. We call an allocation balanced if the bundles of any two agents differ in size by at most one.
Proposition 5.1.
Any balanced allocation achieves a $\frac{k-1}{k}$-MMS approximation under additive leveled valuations for $k\geq 2$, where $k=\lfloor m/n\rfloor$. This bound is tight for balanced allocations.
Proof.
Let $B=(B_{1},\dots,B_{n})$ be a balanced allocation, and let $B^{\prime}=(B^{\prime}_{1},\dots,B^{\prime}_{n})$ be the MMS allocation of agent $i$. Also, denote the bundle corresponding to the MMS value of agent $i$ by $B_{i}^{\prime}$, i.e., $v_{i}(B^{\prime}_{i})=\mu_{i}$. Since $B$ is balanced, $|B_{i}|$ is either equal to $k$ or $k+1$; in the latter case, $i$ trivially guarantees her MMS. We examine the former case. One can observe that due to leveled valuations, $|B^{\prime}_{i}|=k$. Also, for any bundle $B_{l}$ of size $k$ and for any item $j$ of $B^{\prime}_{i}$, it holds $v_{i}(B_{l})\geq v_{i}(B^{\prime}_{i}\setminus\{j\})$. By summing up all these inequalities, we get $k\cdot v_{i}(B_{l})\geq\sum_{j\in B^{\prime}_{i}}v_{i}(B_{i}^{\prime}\setminus\{j\})=(k-1)\cdot v_{i}(B_{i}^{\prime})$ which implies that $v_{i}(B_{l})\geq(k-1)/k\cdot\mu_{i}$, as needed.
Next we show that this bound is tight for balanced allocations. We show that there are instances in which there exists a balanced allocation that guarantees to some agent $i$ only a $\frac{k-1}{k}$ fraction of her MMS. Consider an instance with $n$ agents and $m=kn+1$ items, for some integer parameter $k\geq 1$. We define identical additive valuations with two types of items, namely, a set $L$ of $k$ small items for which each agent has value equal to $\frac{k-1}{k^{2}}+\epsilon$, for sufficiently small $\epsilon$ and a set $H$ of $k(n-1)+1$ large items for which each agent has value equal to $\frac{1}{k}$. It is easy to verify that the proposed valuations are indeed leveled: for any positive integer $t\leq k$, any $t$ small items are always larger than any $(t-1)$ large ones. The MMS value of each agent is at least $1$: consider the allocation in which $n-1$ agents receive a bundle of $k$ large items, each having a total value of $k\cdot 1/k=1$, and one agent gets $k$ small items and one large item, yielding a total value of $k\cdot(\frac{k-1}{k^{2}}+\epsilon)+\frac{1}{k}=1+k\epsilon$. Now, consider a balanced allocation that assigns to agent $i$ all the small items; $i$ receives approximately a $\frac{k-1}{k}$-fraction of her MMS for small enough $\epsilon$. ∎
The mechanism described in Section 4 falls under the broad class of serial dictatorships with quotas (see Definition 5). Serial dictatorships with quotas satisfy strategyproofness.